# Boundary layers, transport and universal distribution in boundary driven active systems

### v2

@@ Appendix D Distribution of ABPs on a semi-infinite line with absorbing condition at $x=0$
$\dot{x}=v\,\cos\theta(t)+\eta(t),~{}\dot{\theta}=\sqrt{2\,D_{r}}\,\zeta(t),$ (D.1)
where $v$ is the self-propulsion speed, $\theta$ is the instantaneous orientation of the particle which takes on continuous values, and $\eta$ is the thermal noise; here $\theta$ evolves as a Brownian particle with (rotational) diffusivity $D_{r}\,$. Since $\langle\cos(\theta(t))\cos(\theta(t^{\prime}))\rangle=e^{-D_{r}\,|t-t^{\prime}|}$, the dynamics of ABP is driven by an exponentially correlated noise.
We simulated the Eq. (D.1) with $\eta(t)=0$ on a semi-infinite line subject to the absorbing boundary condition at $x=0$. We have taken random initial orientation such that $u_{0}=0$. The distribution at large times is captured remarkably well with $P^{\infty}(x,t)-P_{\rm BL}$ given in Eqs. (72)-(73), with $D_{e}=\frac{v^{2}}{2\,D_{r}},l_{M}\approx 0.8\,\frac{v}{D_{r}}\,$. The result is shown in Figure 10.
Figure 10: Distribution of athermal ABPs on a semi-infinite line with an absorbing barrier at $x=0$. The data is shown with particle parameters $v=\sqrt{2},D_{r}=0.2$ and initial condition $x_{0}=1.0,u_{0}=\langle\cos(\theta(0))\rangle\approx 0$. The line corresponds to the expression in Eq. (72) except the boundary layer contribution.
This corroborates the proposed universality of the large time distribution. Note that $P(x,t)$ has a discontinuity at $x=0$. Since the noise is continuous-valued, we expect a boundary layer to be present that will alter the expression of the distribution near $x=0$. This is however not taken up in this paper.

### v4

@@ Appendix F Distribution of ABPs on a semi-infinite line with absorbing condition at $x=0$
$\dot{x}=v\,\cos\theta(t)+\eta(t),~\dot{\theta}=\sqrt{2\,D_{r}}\,\zeta(t),$ (F.1)
where $v$ is the self-propulsion speed, $\theta$ is the instantaneous orientation of the particle which takes on continuous values, $D_{r}$ is rotational diffusivity, and $\eta,\zeta$ are Gaussian white noises. Since $\langle\cos(\theta(t))\cos(\theta(t^{\prime}))\rangle=e^{-D_{r}\,|t-t^{\prime}|}$, the ABP is driven by an exponentially correlated noise.
We simulated the Eq. (F.1) with $\eta(t)=0$ on a semi-infinite line subject to the absorbing boundary condition at $x=0$. We have taken random initial orientation such that $u_{0}=0$. The distribution at large times is captured remarkably well with $P^{\infty}(x,t)-P_{\rm BL}$ given in Eq. (68), with $D_{e}=\frac{v^{2}}{2\,D_{r}},l_{M}\approx 0.8\,\frac{v}{D_{r}}\,$. The result is shown in Figure F.1.
Figure F.1: Distribution of athermal ABPs at different times on a semi-infinite line with an absorbing barrier at $x=0$. The data is shown with particle parameters $v=\sqrt{2},D_{r}=0.2$ and initial condition $x_{0}=1.0,u_{0}=\langle\cos(\theta(0))\rangle\approx 0$. The lines correspond to the expression in Eq. (68).
This corroborates the proposed universality of the large time distribution. Note that $P(x,t)$ has a discontinuity at $x=0$. Since the noise is continuous-valued, we expect a boundary layer to be present near $x=0$. This is however not taken up in this paper.